Fermi National Laboratory
 

Operation Summary for January 5, 2011 - January 7, 2011

 

Wednesday Jan. 5

The day shift began with the Tevatron (TEV) sector B1 cryo system in cooldown mode, with the TEV A-E enclosures and Transfer Hall in access, with the Linac, Booster, and Main Injector (MI) operating normally, and with NuMI and MiniBooNE taking beam.

During the day shift, Operations had the Antiproton Source (Pbar) in standby for work on the Debuncher Lambertson (D:ELAM) power supply, which wrapped up at 11 AM. Operations prepared the Pbar rings for a one-hour access at 10 AM as part of Pbar shielding studies. Pbar resumed stacking at 11:05 AM. The TEV safety system testing ended at 3:30 PM. Operators began a TEV hipot.

During the evening shift, Operations reported that the cooldown of the TEV B1 cryo system had begun and was cold by 11:40 PM. Operations received permission to power the TEV. Operators switched on feeder 23 and started the TEV turn on and checkout. Operations reported that two "network storms" occurred and caused about three hours of downtime. Several systems caused problems after the storm: the Time Line Generator (TLG) needed a reboot; the Booster Gradient Magnet Power Supply (GMPS) crate needed to be power cycled; Booster beam positions were bad and Proton Source personnel tuned for most of the shift; communication with the Pelletron control system (Accelnet) was lost and Pelletron personnel needed to reset the I/O card in the LINUX box; and, a Pelletron blower motor tripped off and required an access to reset.

Thursday Jan. 6

The midnight shift began with the TEV turn on in progress, with Operations monitoring the Recycler stash (115.7E10), with the Antiproton Source (Pbar) stacking (86.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported having current on the TEV bus at 2:40 AM. Operations reported problems with a TEV RF station (TRF8), with the E0 downstream beam valve, and with a low beta magnet quadrupole power supply (B0Q6). These problems will be investigated on the day shift.

During the day shift, an access was required to melt an iceball on the E0 downstream beam valve, this was completed by 10:57 AM. RF technicians repaired TRF8, the work was completed by 11:25 AM. The Tevatron was ramping and Tevatron experts began studies at 12:38 PM. The Tev quenched at 1:49 PM and was recovered by 3:30 PM. Meson Test began taking beam at 3:38 PM.

During the evening shift, Tevatron studies continued. Meson Test was turned off at 6:04 PM. The Tevatron was put into shot setup at 10:13 PM and store 8395 was colliding at 11:21 PM with an Average Initial luminosity (AIL) of 262.1E30.

Friday Jan. 7

The midnight shift began with Operations monitoring store 8395 (200.47 E30) and the stash (274.9 E10), with Pbar stacking (33.5 E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Meson Test began taking beam at 4:19 AM. Store 8395 was terminated at 7:30 AM so that D0 could make an aceess to replace a bad power supply.

The Plans for Friday and the Weekend

The plan for today is to put in a store after the bad D0 power supply has been replaced. Then stack, stash and store over the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/07/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory